OpenStudy (anonymous):

Solve the compound inequality |3x-9|≤15 and |2x-3|≥5. Give answer in interval notation.

OpenStudy (anonymous):

a 3x-9<=15 b 3x-9>=-15 c 2x-3>=5 d 2x-3<=-5 a 3x<= 24 x <=8 b 3x >= -6 x >= -2 [-2,8] c 2x >= 8 x >= 4 d 2x <= -2 x <= -1 [4,infinity) (-infinity,-1) [-2,8] and ( [4,infinity) or (-infinity,-1] ) answer: [-2,-1] or [4,8]

jhonyy9 (jhonyy9):

3x-9<=15 3x<=24 x<=24/3 x<= 8 3x-9<= -15 3x<= -6 x<= -6/3 x<= -2

OpenStudy (anonymous):

what does it mean by compound inequality..?

OpenStudy (anonymous):

"A compound inequality is two or more inequalities joined together with or or and.."

OpenStudy (anonymous):

why is it (-2,-1) or (4,8) instead of (8,4)?

OpenStudy (anonymous):

(low value,high value)

OpenStudy (anonymous):

is there a reason for that?

OpenStudy (anonymous):

i guess because numbers are from left to right..

OpenStudy (anonymous):

ahh. ok thanks ^_^

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

still online..?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

why are the 44 and 1 with infinite but not the 8 and 2?

OpenStudy (anonymous):

i mean 4

OpenStudy (anonymous):

a 3x<= 24 x <=8 b 3x >= -6 x >= -2 looks like -2 <= x <= 8 you see its bounded .. c 2x >= 8 x >= 4 d 2x <= -2 x <= -1 x>=4 but is not smaller than anyvalue so .. 4,infinity the same with the -1

OpenStudy (anonymous):

oh i see